# P2 Explore: Efficient Exploration in Unknown Cluttered Environment with Floor Plan Prediction

### v2

@@ I Introduction
Therefore, a question arises: in the exploration task, is it possible to find the global shortest trajectory? Inspired by the map coverage problem [17], for a fully known environment, we can find an optimal trajectory to achieve coverage of the environment. The primary difference between exploration and map coverage lies in whether prior information about the environment is available. Therefore, if we can predict the status of the unobserved parts based on a partially observed environment, it would be possible to generate an optimal trajectory for exploration.
Previous works have explored how to achieve map prediction in indoor environments [18, 19, 8]. Indoor environments can be categorized into two types: cluttered environments and floor plans. Floor plans are the structure of walls, while cluttered environments are the combination of walls and various objects.
@@
In conclusion, there are two major challenges in using map prediction to accelerate exploration. Firstly, how to achieve efficient map prediction in general cluttered environments. Secondly, how to use inaccurate prediction to accelerate exploration.
In our work, we minimize the gap of current research in predicting maps for cluttered scenarios of any size and how to implement them for exploration robustly. We propose a framework called P2 Explore (Floor Plan Prediction Explore). For the first challenge, rather than directly predicting the scene with obstacles, we focus on predicting the floor plans that are more feasible to achieve. For the second one, we extract the segmentation of rooms in the predicted map and use their connectivity to provide high-level guidance for exploration. To be more specific, firstly, we propose a new architecture of UNet for floor plan prediction called FPUNet. It is trained on a dataset containing real-world scenes. Then, we combine predictions from different frontiers, enabling the generation of a global predicted map. As a result, our algorithm can be extended to maps of any scale. Furthermore, we perform room segmentation based on the predicted map, extracting the specific locations of individual rooms and automatically generating the topological connectivity between them. Then, we optimize the visiting order of all rooms including the predicted ones to accelerate exploration. The illustration for the proposed method can be found in Fig. 2. Our main contributions are:
  • FPUNet is proposed for map prediction in cluttered environment, which is the state-of-the-art (SOTA) archtechture for this task.
@@ III-A Comparison of Map Prediction Results
Figure 4: Illustration for map prediction in different scenes. The first picture in each group is the predicted floor plan. The second picture is the ground truth scene. The scene is getting larger from (a) to (d).
@@ III-B Exploration Results
We compare P2 Explore with three different exploration strategies. The first one is NBV [14], where the next goal for the robot is selected as the frontier with maximum utility. The second one is TSP[5, 13], which can be seen as the SOTA traditional method. In TSP, the robot considers the distribution of all frontiers during each step and formulates a Traveling Salesman Problem (TSP) to traverse all the frontiers. These two methods are algorithms based on current observation information. We also compare P2 Explore with exploration strategies based on predictions [18]. However, in [18], the VAE-based map predictor performed poorly in cluttered environments. Therefore, we use the FPUNet proposed in this paper as its map predictor. [18] acquires predicted information gain based on the predicted map, and selects the next exploration frontier point based on the predicted information gain and path cost.
@@
Method Small ($<$200 m2) Middle (200-600 m2) Large ($>$600 m2)
NBV 87.20 (42.54) 379.55 (110.12) 948.34 (486.97)
TSP 71.77 (38.85) 261.65 (94.18) 652.31 (402.99)
[18] 62.50 (27.76) 255.08 (81.76) 654.76 (339.06)
@@ IV Conclusion & Future Work
In this work, we focus on facilitating the efficiency of exploration using map prediction in cluttered environments. Firstly, we perform floor plan prediction using FPUNet which is the SOTA method for this task. Additionally, we also extract the positions of rooms and their connectivity based on the predicted results. The optimized visiting order of rooms is used to provide a high-level guidance for exploration. The effectiveness of the proposed method is demonstrated by exploration in unknown environments. Our method can shorten the path length compared with the various baselines.
Currently, the algorithm has not been deployed in a real robot. In future work, more comprehensive experimental validation can be conducted to verify its feasibility in a real-world setting.

### v3

@@ I Introduction
Therefore, a question arises: in the exploration task, is it possible to find the global shortest trajectory? Inspired by the map coverage problem [17], for a fully known environment, we can find an optimal trajectory to achieve coverage of the environment. The primary difference between exploration and map coverage lies in whether prior information about the environment is available. Therefore, if we can predict the status of the unobserved parts based on a partially observed environment, it would be possible to generate an optimal trajectory for exploration. Previous works have explored how to achieve map prediction in indoor environments [18, 19, 8] which can be categorized into two types: cluttered environments and floor plans. Floor plans are the structure of walls, while cluttered environments are the combination of walls and various objects.
@@
In conclusion, there are two major challenges in using map prediction to accelerate exploration. Firstly, how to achieve efficient map prediction in general cluttered environments. Secondly, how to use inaccurate prediction to accelerate exploration. In our work, we minimize the gap of current research in predicting maps for cluttered scenarios of any size and how to implement them for exploration robustly. We propose a framework called P2 Explore (Floor Plan Prediction Explore). For the first challenge, rather than directly predicting the scene with obstacles, we focus on predicting the floor plans that are more feasible to achieve. For the second one, we extract the segmentation of rooms in the predicted map and use their connectivity to provide high-level guidance for exploration. To be more specific, firstly, we propose a new architecture of UNet for floor plan prediction called FPUNet. It is trained on a dataset containing real-world scenes. Then, we combine predictions from different frontiers, enabling the generation of a global predicted map. As a result, our algorithm can be extended to maps of any scale. Furthermore, we perform room segmentation based on the predicted map, extracting the specific locations of individual rooms and automatically generating the topological connectivity between them. Then, we optimize the visiting order of all rooms including the predicted ones to accelerate exploration. The illustration for the proposed method can be found in Fig. 2. Our main contributions are:
  • FPUNet is proposed for map prediction in cluttered environment, which is the state-of-the-art (SOTA) architecture.
@@ III-A Comparison of Map Prediction Results
Figure 4: Illustration for map prediction in different scenes. Grid map: observations; Masks: predicted maps. The first picture in each group is the predicted floor plan. The second picture is the ground truth scene. The scene is getting larger from (a) to (d).
@@ III-B Exploration Results
We compare P2 Explore with three different exploration strategies. The first one is NBV [14], where the next goal is selected as the frontier with maximum utility. The second one is TSP[5, 13], which can be seen as the SOTA traditional method. In TSP, the robot considers the distribution of all frontiers during each step and uses TSP algorithm to traverse all the frontiers. These two methods are algorithms based on current observation information. We also compare P2 Explore with strategies based on predictions [18]. However, in Pre-based [18], the VAE-based map predictor performed poorly in cluttered environments. Therefore, we use the FPUNet proposed in this paper as its map predictor. [18] acquires predicted information gain based on the predicted map, and selects the next exploration frontier point based on the predicted information gain and path cost.
@@
Method
Small
($<$200 m2)
Middle
(200-600 m2)
Large
($>$600 m2)
NBV [14] 87.20 (42.54) 379.55 (110.12) 948.34 (486.97)
TSP [5] 71.77 (38.85) 261.65 (94.18) 652.31 (402.99)
Pre-based[18] 62.50 (27.76) 255.08 (81.76) 654.76 (339.06)
@@ IV Conclusion & Future Work
In this work, we focus on facilitating the efficiency of exploration using map prediction in cluttered environments. Firstly, we perform floor plan prediction using FPUNet which is the SOTA method for this task. Additionally, we also extract the positions of rooms and their connectivity based on the predicted results. The optimized visiting order of rooms is used to provide a high-level guidance for exploration. The effectiveness of the proposed method is demonstrated by exploration in unknown environments. Our method can shorten the path length compared with the various baselines. Currently, the algorithm has not been deployed in a real robot. In future work, more comprehensive experimental validation can be conducted to verify its feasibility in a real-world setting. Besides, lightweight network architectures can be considered to enable map prediction.